OpenStudy (4meisu):

You are going to invest $25,000, part at 14% and part at 16%. What is the most that can be invested at 14% to make at least $3600 interest per year? Help please, this has to be an inequality problem.

OpenStudy (anonymous):

.16x + .14*(25000 - x) = 3600 .16x + 3500 - .14x = 3600 .02x = 100 x = 100/.02 =5000 this is for the 16 % to find investment at 14% $25000 - $5000 = $20000 if you say that x= the amount invested at 14% you have this .14x + .16*(25000-x) =3600 .14x + 4000 - .16x =3600 -.02x = -400 x = -400/-.02 =$20000